AS8051(1L) AS8051(1L) NAME as8051−crossassemblerformicrocomputers SYNOPSIS as8051[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input DESCRIPTION Theas8051commandassemblestheinputfileintoatextoutputfilerepresentingtheprogram memoryforamicrocomputer. Options -pcpuIgnoredbythisprogram. -llistfile Outputafileformatedwiththememoryaddress,data,andsourceinputlines. -ohexfile Outputthememorydataimageinaformacceptedbymostpromprogrammers. -hhexfile Sameas−o. −dSavetheintermediatedatafile(seeFILE)andaborttheexecutionattheendofprocessing. -ssymbolfile Printthesymboltablevaluesandnames,oneperlineinthespecifiedfile. FILES /usr/tmp/frtXXXXXX SEEALSO FrankensteinCrossAssemblersUsersManual(filebase.doc)Appendixforas8051Frankenstein Assembler(fileas8051.doc) NOTES Thereisonlyoneinputfile. 1